# TimeDistill: Efficient Long-Term Time Series Forecasting with MLP via Cross-Architecture Distillation

### v1

@@ Appendix I Efficiency Analysis
Figure 12: Comparison of model predictions across temporal scales and corresponding spectrograms before and after distillation on the ETTm1 dataset. MSE values for MLP, Teacher (ModernTCN (Luo & Wang, 2024)), and TimeDistill are 2.16, 0.73 and 0.74.
Figure 13: Comparison of model predictions across temporal scales and corresponding spectrograms before and after distillation on the ETTm2 dataset. MSE values for MLP, Teacher (ModernTCN (Luo & Wang, 2024)), and TimeDistill are 3.35, 1.19, 1.28.
Appendix L Showcases
@@
For clear comparison, we present test set showcases in Appendix P, where TimeDistill shows better performance.
Appendix M Full Main Results
@@
Table 12: Long-term time series forecasting results with prediction lengths $S~{}\in\{96,192,336,720\}$. A lower MSE or MAE indicates a better prediction. For consistency, we maintain a fixed input length of 720 throughout all the experiments. The best performance is highlighted in red, and the second-best is underlined.
@@
(Ours) (2024) (2024) (2024a) (2023) (2023) (2022) (2022) (2021)
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 14: Model efficiency comparison averaged across all prediction lengths (96, 192, 336, 720) on eight datasets. The size of each bubble represents the number of parameters in the corresponding method, with larger bubbles indicating more parameters.
Appendix N Full Different Look-back Window Results
Table 13: Results with look-back length $S=96$.
@@
Table 14: Results with look-back length $S=192$.
@@
Table 15: Results with look-back length $S=336$.
@@
Table 16: Results with look-back length $S=720$.
@@
Appendix O Full Ablation Results
Table 17: Ablation study on different datasets (Teacher: iTransformer (Liu et al., 2024)).
@@
Table 18: Ablation study on different datasets (Teacher: ModernTCN (Luo & Wang, 2024)).
@@
Table 19: Ablation study measured by MSE on different components of TimeDistill. Teacher is iTransformer;
@@
Table 20: Ablation study measured by MSE on different components of TimeDistill. Teacher is ModernTCN.
@@
Appendix P Show Cases
Figure 15: Prediction cases from ECL by different models under the input-720-predict-96 settings. Black lines are the ground truths and Red lines are the model predictions. Due to space constraints, we only retained the last 96 time steps of input for plotting.
Figure 16: Prediction cases from ETTh1 by different models under the input-720-predict-96 settings.
Figure 17: Prediction cases from ETTh2 by different models under the input-720-predict-96 settings.
Figure 18: Prediction cases from ETTm1 by different models under the input-720-predict-96 settings.
Figure 19: Prediction cases from ETTm2 by different models under the input-720-predict-96 settings.
Figure 20: Prediction cases from Traffic by different models under the input-720-predict-96 settings.
Figure 21: Prediction cases from Solar by different models under the input-720-predict-96 settings.
Figure 22: Prediction cases from Weather by different models under the input-720-predict-96 settings.

### v2

@@ Appendix I Directly Match the Patterns in Ground Truth Time Series
Table 12. Comparison of SpectralEntropy and Variance Ratios for different datasets
Dataset SpectralEntropy VarRatio1H VarRatio6H VarRatio1D
ECL 7.77 1 0.84 0.76
ETTh1 6.38 1 0.71 0.23
ETTm1 6.46 1 0.70 0.23
Traffic 7.45 1 0.55 0.07
Linear 2.36 1 1 1
Appendix L What if there is no Periodic or Multi-Scale Structure in the Dataset?
TimeDistill may perform suboptimally on time series without strong periodicity or multi-scale patterns. However, such cases are rare, as most real-world time series exhibit these properties (timesnet; autoformer). To verify this, we analyze four datasets using spectral entropy and variance ratio (1H/6H/1D denote different scales). Spectral entropy measures the diversity of periodic components, while variance ratio captures changes across scales. High entropy and decreasing variance ratios indicate strong periodic and multi-scale structure. As shown in the Table 12, real-world datasets show both, unlike synthetic linear series, supporting TimeDistill’s broad applicability.
Moreover, TimeDistill does not depend solely on multi-period distillation. As shown in Table 4 in the paper, even without it, the framework still improves student performance. Additionally, the loss weights are adjustable, allowing adaptation to datasets with weak periodicity.
Furthermore, Theorem 4.1 and Theorem 4.2 show that the multi-scale and multi-period loss act as mixup data augmentation between the teacher’s and ground truth period distributions. This augmentation remains useful even with weak multi-scale or periodic structure, as long as a distributional gap exists.
Table 13. Comparison with other KD methods tailored for time series forecasting. Avg.Imp. denotes the average relative improvement over MLP.
Dataset MLP TimeDistill DE-TSMCL LightCTS*
ETTh1 0.502 0.429 0.479 0.448
ETTh2 0.393 0.345 0.376 0.389
ETTm1 0.391 0.348 0.386 0.373
ETTm2 0.300 0.244 0.268 0.271
Weather 0.234 0.220 0.223 0.224
Avg.Imp. - 12.41% 5.17% 5.99%
Table 14. Performance comparison on PEMS datasets (PEMS03, PEMS04, PEMS07, PEMS08, and average).
Dataset Metric MLP TimeMixer TimeDistill-TimeMixer
PEMS03 MAE 16.33 15.73 14.86
MAPE 16.78 15.89 14.14
RMSE 26.19 25.85 24.03
PEMS04 MAE 22.47 20.00 19.96
MAPE 15.34 13.94 12.26
RMSE 35.39 32.80 32.65
PEMS07 MAE 22.64 20.88 20.30
MAPE 10.03 9.88 8.40
RMSE 36.16 33.57 33.48
PEMS08 MAE 17.22 14.89 15.02
MAPE 11.07 9.39 9.29
RMSE 27.83 24.00 24.92
Average MAE 19.67 17.88 17.53
MAPE 13.31 12.28 11.02
RMSE 31.39 29.06 28.77
Appendix M Comparison with KD Tailored for Time Series Forecasting
[…]
Appendix N Short-term Time Series Forecasting Performance
TimeDistill can be directly adapted to the short-term time series forecasting task. Following TimeMixer (timemixer), we provide some preliminary results in Table 14. We report the results over four PEMS datasets. TimeDistill significantly improves MLP’s performance and surpasses its teacher on the short-term forecasting task.
Appendix O Future Directions
Distilling from advanced time–series models.
Section 5.3 shows that TimeDistill can already learn from a wide spectrum of teachers—ModernTCN, iTransformer, PatchTST, and TimeMixer—yielding MSE reductions of up to $13.9\%$ for the student and up to $10.4\%$ over the teachers themselves (Table F). The conclusion, therefore, proposes “distilling from advanced time-series (foundation) models” as the next milestone. Concrete directions include: (i) treating recently released foundation checkpoints (e.g., TimesFM) as teachers and analysing how their broad temporal representations propagate through the multi-scale/multi-period objectives in Figure 6; and (ii) leveraging heterogeneous teacher ensembles (Appendix F, Table 9) to let a single student interpolate knowledge from climate, finance, and health domains.
Incorporating multivariate patterns explicitly.
Although the current student is channel–independent, Section 5.4 reveals that distillation already aligns its inter-variable correlation matrix with that of a channel-dependent teacher (Figure 9). Future work can (a) add a covariance-matching loss to reinforce this implicit transfer, and (b) draw on teachers specialised in cross-channel modelling, such as iTransformer. A realistic test-bed is the Traffic dataset, where TimeDistill still lowers MSE by $10.3\%$ despite 861 variables (Table F).
Extending the framework to imputation and classification.
Knowledge distillation is task-agnostic: Appendix N confirms that the same student–teacher pair handles short-term forecasting. Because the multi-scale and multi-period losses operate on reconstructed values and spectra, replacing the forecasting objective with a masked reconstruction loss would adapt TimeDistill to missing-value imputation. Moreover, the mixup-based view in Theorem 4.1 naturally generalises the framework to classification by softening class targets. Validating these ideas on UCR/UEA (classification) and M4C (imputation) benchmarks—while re-using the qualitative analyses of Figure 8 to track temporal fidelity—is an immediate next step.
Table 15. Comparison of methods on different datasets under three settings: Original, Non-stationary, and REVIN.
None Non-stationary REVIN
Dataset TimeDistill iTransformer ModernTCN TimeMixer TimeDistill iTransformer ModernTCN TimeMixer TimeDistill iTransformer ModernTCN TimeMixer
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ECL 0.160 0.259 0.166 0.267 0.166 0.268 0.174 0.270 0.157 0.254 0.163 0.259 0.166 0.268 0.162 0.257 0.161 0.255 0.160 0.257 0.159 0.256 0.165 0.259
ETTh1 0.443 0.450 0.514 0.521 0.508 0.507 0.460 0.472 0.429 0.441 0.468 0.476 0.508 0.507 0.440 0.451 0.433 0.448 0.468 0.477 0.443 0.447 0.459 0.465
ETTh2 0.452 0.460 1.209 0.826 0.783 0.606 0.862 0.645 0.345 0.395 0.398 0.426 0.783 0.606 0.382 0.412 0.366 0.407 0.406 0.431 0.347 0.396 0.422 0.444
ETTm1 0.362 0.392 0.426 0.453 0.535 0.508 0.374 0.397 0.348 0.379 0.372 0.402 0.535 0.508 0.392 0.409 0.354 0.391 0.374 0.403 0.365 0.388 0.367 0.388
ETTm2 0.258 0.332 0.371 0.410 1.641 0.953 0.322 0.370 0.244 0.311 0.276 0.337 1.641 0.953 0.281 0.332 0.270 0.324 0.275 0.336 0.267 0.330 0.279 0.339
Solar 0.184 0.242 0.207 0.264 0.192 0.246 0.238 0.288 0.184 0.241 0.214 0.270 0.192 0.246 0.238 0.288 0.193 0.248 0.203 0.273 0.191 0.259 0.238 0.288
Traffic 0.395 0.277 0.661 0.407 0.414 0.289 0.426 0.282 0.404 0.278 0.379 0.271 0.414 0.289 0.414 0.289 0.387 0.271 0.385 0.274 0.413 0.284 0.391 0.275
Weather 0.236 0.299 0.301 0.345 0.244 0.310 0.235 0.287 0.220 0.269 0.259 0.290 0.244 0.310 0.232 0.273 0.226 0.270 0.260 0.291 0.229 0.269 0.230 0.271
Average 0.311 0.339 0.482 0.437 0.560 0.461 0.386 0.376 0.292 0.321 0.316 0.341 0.560 0.461 0.318 0.339 0.299 0.327 0.316 0.343 0.302 0.329 0.319 0.341
Appendix P Effect of Normalization on Performance Across Methods
Our experiment setup follows standard practice (itransformer; non-stationary; sparsetsf) to ensure fairness: each model uses its best-performing normalization based on validation. To reduce confounding factors, we also evaluate all models under the same normalization. As shown in Tables 15 (Non-stationary and REVIN), TimeDistill consistently outperforms prior methods, confirming that its gains come from the distillation method rather than preprocessing differences.
We also study the impact of normalization by removing it from both TimeDistill and baselines in Table 15 (None). Results show that TimeDistill consistently outperforms all baselines. This confirms that its gains are not due to preprocessing, but to the method itself, improving interpretability.
Appendix Q Showcases
@@
For clear comparison, we present test set showcases in Appendix U, where TimeDistill shows better performance.
Figure 12. Comparison of model predictions across temporal scales and corresponding spectrograms before and after distillation on the ETTm1 dataset. MSE values for MLP, Teacher (ModernTCN (moderntcn)), and TimeDistill are 2.16, 0.73, and 0.74.
Figure 13. Comparison of model predictions across temporal scales and corresponding spectrograms before and after distillation on the ETTm2 dataset. MSE values for MLP, Teacher (ModernTCN (moderntcn)), and TimeDistill are 3.35, 1.19, and 1.28.
Table 16. Performance promotion obtained by our TimeDistill framework with different student models. The notation nL-d represents the structure of an MLP model, where nL specifies the number of layers and d denotes the dimension of the hidden units in each layer. For instance, 2L-512 refers to a model with 2 layers and a hidden dimension of 512. We report the average performance from all prediction lengths. $\Delta_{Student}$ represents the performance promotion between Student+TimeDistill and original trained student.
Student Models MLP MLP MLP MLP LightTS TSMixer FITS
(2L-512) (3L-512) (4L-512) (2L-1024)
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 Student 0.502 0.489 0.487 0.497 0.468 0.481 0.495 0.500 0.465 0.471 0.471 0.474 0.428 0.443
+TimeDistill 0.428 0.445 0.442 0.448 0.443 0.453 0.442 0.446 0.436 0.445 0.433 0.446 0.411 0.430
$\Delta_{Student}$ 14.74% 9.00% 9.11% 9.86% 5.29% 5.70% 10.79% 10.83% 6.26% 5.57% 8.02% 5.92% 3.96% 2.75%
ETTh2 Student 0.393 0.438 0.742 0.601 0.908 0.668 0.639 0.563 0.675 0.582 0.376 0.420 0.338 0.388
+TimeDistill 0.345 0.397 0.341 0.390 0.362 0.406 0.344 0.390 0.382 0.423 0.351 0.399 0.333 0.382
$\Delta_{Student}$ 12.16% 9.46% 53.97% 35.05% 60.12% 39.19% 46.20% 30.72% 43.42% 27.26% 6.83% 5.01% 1.52% 1.70%
ETTm1 Student 0.391 0.413 0.378 0.406 0.392 0.414 0.377 0.402 0.376 0.399 0.370 0.395 0.357 0.379
+TimeDistill 0.354 0.390 0.346 0.380 0.353 0.389 0.346 0.380 0.358 0.382 0.354 0.383 0.358 0.380
$\Delta_{Student}$ 9.42% 5.65% 8.54% 6.34% 9.90% 6.12% 8.40% 5.56% 4.83% 4.29% 4.13% 2.92% -0.45% -0.27%
ETTm2 Student 0.300 0.373 0.359 0.408 0.346 0.393 0.307 0.374 0.283 0.346 0.295 0.347 0.253 0.316
+TimeDistill 0.252 0.316 0.259 0.321 0.254 0.317 0.245 0.312 0.258 0.324 0.253 0.319 0.248 0.311
$\Delta_{Student}$ 16.07% 15.28% 27.77% 21.31% 26.71% 19.41% 20.03% 16.55% 8.92% 6.38% 14.19% 8.23% 2.20% 1.40%
Weather Student 0.234 0.294 0.232 0.289 0.237 0.291 0.221 0.278 0.235 0.293 0.239 0.278 0.241 0.281
+TimeDistill 0.220 0.270 0.222 0.267 0.225 0.271 0.220 0.267 0.221 0.272 0.222 0.266 0.239 0.279
$\Delta_{Student}$ 5.83% 8.02% 4.53% 7.53% 5.22% 7.04% 0.53% 4.10% 5.95% 7.17% 7.19% 4.09% 0.90% 0.74%
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 14. Model efficiency comparison averaged across all prediction lengths (96, 192, 336, 720) on eight datasets. The size of each bubble represents the number of parameters in the corresponding method, with larger bubbles indicating more parameters.
Appendix R Full Main Results
@@
Table 17. Long-term time series forecasting results with prediction lengths $S~\in\{96,192,336,720\}$. A lower MSE or MAE indicates a better prediction. For consistency, we maintain a fixed input length of 720 throughout all the experiments. The best performance is highlighted in red, and the second-best is underlined.
@@
(Ours) ((itransformer)) ((moderntcn)) ((timemixer)) ((patchtst)) ((micn)) ((fedformer)) ((timesnet)) ((autoformer))
@@
Appendix S Full Different Look-back Window Results
Table 18. Results with look-back length $S=96$.
@@
Table 19. Results with look-back length $S=192$.
@@
Table 20. Results with look-back length $S=336$.
@@
Table 21. Results with look-back length $S=720$.
@@
Discussion on although ModernTCN and iTransformer differ in performance, their distilled student models perform very consistently.
There is an interesting observation from the above four tables that although ModernTCN and iTransformer have different performances, the two student models distilled from them as teacher models perform highly consistently. The differences are clear on Traffic, ETTh1/2, Solar, and Weather, but less so on ECL and ETTm1/m2, where students distilled from ModernTCN and iTransformer perform similarly. We attribute this to the limited capacity of the student MLP, leading to performance saturation. Table 22 supports this: as student size increases, performance differences on ETTm1 become more apparent, confirming our hypothesis.
Table 22. Performance comparison (MSE and MAE) for different MLP sizes on ModernTCN and iTransformer.
Method MLP(2L512)MSE MLP(2L512)MAE MLP(2L1024)MSE MLP(2L1024)MAE MLP(2L2048)MSE MLP(2L2048)MAE
ModernTCN 0.376 0.395 0.374 0.393 0.371 0.382
iTransformer 0.376 0.395 0.376 0.394 0.374 0.389
Appendix T Full Ablation Results
Table 23. Ablation study on different datasets (Teacher: iTransformer (itransformer)).
@@
Table 24. Ablation study on different datasets (Teacher: ModernTCN (moderntcn)).
@@
Table 25. Ablation study measured by MSE on different components of TimeDistill. Teacher is iTransformer;
@@
Table 26. Ablation study measured by MSE on different components of TimeDistill. Teacher is ModernTCN.
@@
Discussion on removing multi-period loss can still yield acceptable results.
We observe that removing multi-period distillation causes a small drop when teacher is ModernTCN, but a large drop with iTransformer (Table 23 and Table 25).
Theorem 4.2 shows that multi-period loss optimizes a mixup distribution $\mathbf{Q}_{y}+\lambda\,\mathbf{Q}_{t}$. When $\mathbf{Q}_{t}$ is already close to $\mathbf{Q}_{y}$, the mixup offers limited gain. We confirm this on ETTh1 by computing FFT-based period distributions and KL divergence. As shown in Table 27, ModernTCN’s $\mathbf{Q}_{t}$ is closer to $\mathbf{Q}_{y}$ than iTransformer’s, supporting the hypothesis.
Table 27. Comparison of KL values
iTransformer ModernTCN
KL 0.4606 0.3843
Appendix U Show Cases
Figure 15. Prediction cases from ECL by different models under the input-720-predict-96 settings. Black lines are the ground truths and Red lines are the model predictions. Due to space constraints, we only retained the last 96 time steps of input for plotting.
Figure 16. Prediction cases from ETTh1 by different models under the input-720-predict-96 settings.
Figure 17. Prediction cases from ETTh2 by different models under the input-720-predict-96 settings.
Figure 18. Prediction cases from ETTm1 by different models under the input-720-predict-96 settings.
Figure 19. Prediction cases from ETTm2 by different models under the input-720-predict-96 settings.
Figure 20. Prediction cases from Traffic by different models under the input-720-predict-96 settings.
Figure 21. Prediction cases from Solar by different models under the input-720-predict-96 settings.
Figure 22. Prediction cases from Weather by different models under the input-720-predict-96 settings.